OpenStudy (anonymous):

Estimate the area of the circle. Use 22 over 7

OpenStudy (anonymous):

help?

OpenStudy (shamil98):

area = pir^2 d = 28 r = 14 plug it in

OpenStudy (anonymous):

hm 1176, and then i used a diffrent way and got another asnwer

OpenStudy (anonymous):

0.5 is the other i found what is the right one ?

OpenStudy (math&ing001):

@AngelicMelody check your answers again. Use the formula Shamil gave above, and instead of using π use 22/7 as they asked you to, it should look like this : \[Area = \frac{ r² * 22 }{ 7 }\] with r=d/2=14

OpenStudy (anonymous):

22/7 x 14 = 44

OpenStudy (math&ing001):

You forgot the power 2 and the unit at the end. 22/7 x 14² = 616 cm² Good luck in your studies !